# Will LLMs Scaling Hit the Wall? Breaking Barriers via Distributed Resources on Massive Edge Devices

### v1

@@ Abstract
The remarkable success of foundation models has been driven by scaling laws, demonstrating that model performance improves predictably with increased training data and model size. However, this scaling trajectory faces two critical challenges: the depletion of high-quality public data, and the prohibitive computational power required for larger models, which have been monopolized by tech giants. These two bottlenecks pose significant obstacles to the further development of AI. In this position paper, we argue that leveraging massive distributed edge devices can break through these barriers. We reveal the vast untapped potential of data and computational resources on massive edge devices, and review recent technical advancements in distributed/federated learning that make this new paradigm viable. Our analysis suggests that by collaborating on edge devices, everyone can participate in training large language models with small edge devices. This paradigm shift towards distributed training on edge has the potential to democratize AI development and foster a more inclusive AI community.
Machine Learning, ICML
@@ 1 Introduction
Scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) have been fundamental to the remarkable success of foundation models, demonstrating a predictable relationship between performance and the expansion of model parameters and training data. These laws have guided the development of increasingly powerful models, from BERT (Devlin et al., 2018) to GPT-4 (OpenAI, 2023b), showing that performance improvements can be achieved through systematic scaling of both model size and training data (Brown et al., 2020; Chowdhery et al., 2022). However, the continued application of these scaling laws requires ever-increasing amounts of data and computational resources, pushing the boundaries of what is currently feasible (Patterson et al., 2021).
@@
Public data has been the primary fuel driving AI development forward. This field has witnessed an exponential growth in data requirements, from the early success of MNIST (LeCun et al., 1998b) with its 70,000 handwritten digits to ImageNet’s revolutionary impact with 14 million labeled images (Deng et al., 2009). This trajectory has continued with modern large language models (LLMs) like GPT (OpenAI, 2023a), LLaMA (Meta AI, 2023), and DeepSeek (Liu et al., 2024a) series, which are trained on trillions of tokens. Recent evidence that LLaMA 3.1’s smallest model (8B) (Meta, 2024) trained on 15 trillion tokens, outperforms LLaMA 2’s largest model (70B) (Touvron et al., 2023) trained on 2 trillion tokens (despite being 10×10$\times$10 × smaller in model size, the 7×7$\times$7 × increase in training data leads to superior performance), demonstrates the paramount importance of data scaling (Sun et al., 2017; Raffel et al., 2020; DeepLearning.AI, 2024). However, we are witnessing a concerning trend of data depletion, where high-quality public data sources are becoming exhausted (Lee et al., 2021; Biderman et al., 2022). Villalobos et al. argues that human-generated public text data cannot sustain scaling beyond this decade. While recent efforts advocate for training larger models with synthetic data (Chen et al., 2024a), AI-generated content may fail to yield performance improvements (Wenger, 2024), also risks polluting public data sources (Fang et al., 2024). Moreover, stricter data privacy regulations like GDPR (Parliament & of the European Union, 2016) have made data collection increasingly difficult and expensive. This looming data scarcity suggests that scaling laws may hit a wall (Hardy-White, 2024), potentially impeding further AI advancement.
@@
Computational resources has been the primary engine powering AI development. Throughout AI history, major breakthroughs have been closely tied to advances in computing power, from early models requiring single CPUs (with peak performance of 1-2 GFLOPS) to modern GPU clusters. The computational demands have grown exponentially - from BERT-Large’s training requiring 64 TPU v3 chips (providing 420 TFLOPS) (Devlin et al., 2018) to GPT-3’s training on 10,000 V100 GPUs (reaching 28,000 TFLOPS) (Brown et al., 2020), while training GPT-4 reportedly required over 25,000 NVIDIA A100 GPUs (delivering a staggering 400,000 TFLOPS) (OpenAI, 2023b). More recent models like Grok 3 push these requirements even further (xAI, 2025). However, we are approaching physical limits in single-chip performance as Moore’s Law slows down (Thompson et al., 2021). While massive computing clusters can compensate for individual chip limitations, maintaining such infrastructure incurs astronomical costs - estimated at over \$100M for training GPT-4 (Sharir et al., 2020) - and poses significant environmental concerns due to their enormous energy consumption, with each training run emitting as much CO2 as 500 cars driven for a year (Schwartz et al., 2020). Moreover, this level of computing power has become concentrated among a few tech giants, creating a monopolistic landscape that effectively excludes smaller companies and academic institutions from participating in foundational AI research (Ahmed et al., 2022). This centralization of computing resources presents a significant barrier to innovation and democratization in AI development (Thompson et al., 2022).
@@
In this paper, we propose that leveraging massive distributed edge devices offers a promising solution to overcome both data and computing barriers in AI development. Our analysis reveals two compelling opportunities: First, edge data generated from smartphones for past 5 years are projected to reach 33.1 EB, offering fresh, diverse, and contextually rich training samples. Second, the collective computing power of edge devices - with smartphones delivering 9,278 EFLOPS for past 5 years - demonstrates the feasibility of distributed model training, as training state-of-the-art models like DeepSeek-v3 would require only about 60,723 users with edge devices working (ideally) in parallel to match its current training setup. Based on these insights, we argue that leveraging these massive distributed edge devices can break barriers of data and computing wall, and everyone can participate in training large models with small edge devices. To support this position, we first analyze the critical challenges of large language models, examining both data bottlenecks (§2.1) and computational monopolization (§2.2). We then explore the hidden potential of massive edge devices, investigating their vast untapped distributed data resources (§3.1) and computational capabilities (§3.2). Building on these insights, we investigate technical approaches for overcoming large model challenges through distributed computing architectures (§4): small language models at edges (§4.1), collaborative inference (§4.2), and collaborative training (§4.3). We then identify two critical open challenges: heterogeneous device model fusion and heterogeneous device compute sharing (§5). Finally, we discuss the societal impact like AI democratization, incentive mechanisms, and environmental benefits of this paradigm shift (§A).
@@ 2.1 The Ceiling of Public Data
The rapid advancement of large language models has created an insatiable appetite for training data. Neural scaling laws establish that model performance improves predictably with data quantity—a relationship that demands exponentially growing datasets (Hoffmann et al., 2022). A canonical example is GPT-3, trained on 300 billion tokens spanning books, web content, and programming code (Brown et al., 2020). Current projections suggest dataset sizes grow at 0.38 orders of magnitude (2.4$\times$) annually (Villalobos & Ho, 2022), implying models will require three orders of magnitude more data within a decade.
Despite the internet’s vast textual resources, the total stock of high-quality human-generated text remains bounded. Recent estimates place this limit at approximately $4\times 10^{14}$ tokens (Villalobos & Ho, 2022). Villalobos et al. (2024) argues that current consumption patterns suggest exhaustion of public text data by 2028, potentially accelerated to 2026 through excessive data reuse during training (a practice termed overtraining). Therefore, the finite nature of publicly available human-generated text data is expected to become a major bottleneck for LLM scaling within the next decade. Despite the current large scale of public data, the risk of data exhaustion is rapidly approaching as data demand continues to grow (Sevilla et al., 2022).
@@
Faced with the threat of data exhaustion, researchers have proposed various solutions, among which synthetic data generation is considered one of the most promising approaches. By leveraging LLMs to produce their own training data, researchers envision self-sustaining data ecosystems. Early successes in constrained domains like mathematics and code generation, where automated verification ensures quality, demonstrate potential (Liu et al., 2023). Recent work (Chen et al., 2024a) demonstrated that diverse synthetic data enhances the performance of LLMs during both pre-training and fine-tuning.
The adoption of synthetic data faces three fundamental challenges. First, model collapse occurs when models iteratively train on their own outputs, causing gradual divergence from original data distributions. This recursive process amplifies biases and reduces output diversity, ultimately degrading model performance across generations (Shumailov et al., 2023; Dohmatob et al., 2024). Second, synthetic data quality remains inherently unverifiable in open-domain contexts. While formal domains like mathematics allow algorithmic validation, natural language lacks objective evaluation standards. The absence of ground-truth verification creates self-referential quality assessments, compromising reliability (Alemohammad et al., 2023; Wenger, 2024). Finally, synthetic data struggles to replicate human linguistic diversity. Current methods disproportionately replicate dominant language patterns while underrepresenting cultural nuances and low-frequency expressions. This homogeneity limits their utility for training robust general-purpose models (Fan et al., 2023a).
@@ 2.2 The Monopoly of Computing Resources
Figure 1: Trend of Computational Demand for Model Training. (Data source: Notable AI models (Hobbhahn et al., 2023)).
@@
As large-scale AI models like GPT-4 (OpenAI, 2023b), Llama 3 (Meta, 2024), and DeepSeek-V3 (Liu et al., 2024a) surpass the trillion-parameter scale, the global AI landscape faces severe computational efficiency challenges. As shown in Figure 1, since the deep learning revolution in 2010, AI training demands have grown at a super-exponential rate of 3.9$\times$ per year—an acceleration that intensified with the adoption of the transformer architecture as the industry standard (Vaswani et al., 2017). With the advent of the era of large language models in 2022, the demand for computing power has surged even further, reaching an unprecedented growth rate of 13.4$\times$ per year. This marks a transformative shift in AI computation, where the need for computing power is expanding at an unprecedented pace, pushing the limits of existing hardware and infrastructure.
@@
Moore’s Law, which has driven the growth in computing power for decades, is slowing down as we approach the physical limits of silicon-based chip technology (Kressel, 2023). The difficulty in shrinking transistors has led to diminishing returns in computational performance. As a result, the AI industry is relying more on specialized hardware like GPUs, TPUs, and custom chips to meet growing demands. However, this shift has made high-performance hardware even more expensive and exclusive, further intensifying the gap between organizations with the resources to develop advanced AI models and those without.
@@
The rapid expansion of AI model scales and the surge in computational demand are facing dual constraints in global computing infrastructure. On one hand, bottlenecks in advanced semiconductor manufacturing severely limit the expansion rate of AI data centers. The foundry capacity for wafers at 5nm and below—such as those produced by TSMC—has already been fully booked by leading technology companies until 2026 (Staff, 2024). Moreover, the construction of new wafer fabs involves long lead times and is further constrained by the global supply chain shortages of critical equipment, such as lithography machines. On the other hand, the exponential increase in chip deployment within individual AI clusters is putting immense pressure on the already limited semiconductor manufacturing capacity, pushing the industry toward its production ceiling (ScaleFlux Research, 2024). These factors have significantly hindered the continuous expansion of computing power, making it increasingly difficult to scale AI infrastructure sustainably.
@@ 3.1 Massive Data from Edges
Figure 2: Global data volume from 2014 to 2025 and IoT device data volume in 2015 and 2025. (Data sources: Global data volume from (Statista global data volume, 2023); IoT device data volume from (Statista IoT device data volume, 2023).)
Figure 3: Smartphone data volume with edge computing market size (right) from 2018 to 2028. (Data sources: Edge computing market (Grand View Research, 2023); Smartphone data volume derived from the number of smartphones (BankMyCell, 2023).)
As discussed in § 2.1, edge data represents a crucial alternative to synthetic data in addressing the challenge of data exhaustion. Edge data refers to the data generated by edge devices at or near the source of data generation, which typically remains private and localized rather than being publicly accessible. Edge devices encompass a wide range of equipment including Internet of Things (IoT) sensors, smartphones, wearables, industrial controllers, and other smart devices that process data at the network edge. The global data ecosystem exhibits significant bifurcation: publicly available data constitutes merely the tip of the iceberg, while data generated by edge devices demonstrates superior strategic value both in terms of growth rate and high quality.
@@
As illustrated in Figure 3, the global data volume is projected to reach 182 ZB by 2025 (Statista global data volume, 2023), with a particularly pronounced growth in edge-generated data. The data generated by IoT devices is anticipated to increase from 13.6 ZB in 2019 to 79.4 ZB in 2025 (Statista IoT device data volume, 2023), elevating its share of the global data volume from 33.2% to 43.6%. Over the period from 2015 to 2025, the global data volume exhibited a compound annual growth rate (CAGR) of 35.1%, resulting in an overall increase of 1074.2% and a cumulative total of 871.5 ZB. IoT device data experienced a growth of 483.8% from 2019 to 2025. This trend underscores the increasingly central role of edge-generated data in the global data ecosystem. Beyond IoT devices, smartphones, as a critical source of edge-side data, are also contributing to the steady rise in data volume. As depicted in Figure 3, the estimated smartphone data volume is projected to grow from 5 EB in 2018 to 8 EB by 2028  [^1] These numbers are estimated based on an average user data generation of 1 GB per device. For detailed estimation methodology, refer to Appendix C.. This exponential growth is closely aligned with the rapid expansion of the edge computing market, which is forecasted to surge from \$5.5 billion in 2019 to \$87.9 billion by 2028, representing a remarkable growth rate of 3215.1%. The burgeoning edge computing market has further catalyzed the generation and processing of edge-side data, reinforcing its significance in the broader data landscape.
Insight: The smartphone data volume of the past 5 years (before 2025) is projected to reach approximately 33.1 EB, with unique advantages in privacy and real-time context, demonstrating the massive data potential of edge for AI model training.
Edge-generated data is of high quality.
@@
Beyond its impressive quantity, edge data possesses several qualitative advantages that make it particularly valuable for model pretraining. First, edge data exhibits superior diversity across multiple dimensions. It encompasses a wide variety of data types from IoT devices, mobile interactions, and personal devices, covering different domains, languages, and user behaviors. This natural diversity provides richer training signals compared to curated public datasets (Nayak et al., 2024). Second, edge data demonstrates strong real-time capability. Unlike public datasets, which are often updated infrequently, edge devices continuously generate fresh data with low latency (Wireless, 2023), offering more up-to-date and relevant training samples. Third, edge data offers enhanced personalization characteristics. It enables real-time interactions and context-aware recommendations by leveraging local data such as location and device behavior. This allows for highly relevant and adaptive personalization that can dynamically adjust to user preferences (XenonStack, 2023). Fourth, edge data exhibits enhanced context awareness due to the proximal positioning of edge devices to data sources, enabling the capture of fine-grained contextual information often absent in public datasets. For instance, wearable devices collect synchronized activity and location data, providing high-fidelity behavioral patterns that preserve temporal and spatial context.
@@
In conclusion, edge data with its explosive growth 222 Seagate (2019, 2020) provide a more comprehensive overview of the global data volume, but we cannot access the statistics data. We appreciate any suggestions for better statistical data sources. and superior qualitative characteristics, is a valuable resource for model pretraining. Its diversity, real-time nature, personalization, and rich context make it an ideal foundation for developing robust and adaptable large-scale models, enhancing their ability to serve real-world applications effectively.
Figure 4: Edge Computing Power Evolution Trend. (Data source: (NanoReview.net, 2025)).
Figure 5: Smartphone Market Share and Computing Power Trends. (Data source: (Canalys, 2025)).

### v2

@@ Abstract
The remarkable success of foundation models has been driven by scaling laws, demonstrating that model performance improves predictably with increased training data and model size. However, this scaling trajectory faces two critical challenges: the exhaustion of high-quality public data, and the prohibitive computational power required for larger models, which have been monopolized by tech giants. These two bottlenecks pose significant obstacles to the further development of AI. In this position paper, we argue that leveraging massive distributed edge devices can break through these barriers. We reveal the vast untapped potential of data and computational resources on massive edge devices, and review recent technical advancements in distributed/federated learning that make this new paradigm viable. Our analysis suggests that by collaborating on edge devices, everyone can participate in training large language models with small edge devices. This paradigm shift towards distributed training on edge has the potential to democratize AI development and foster a more inclusive AI community. The project page is available at https://tao-shen.github.io/Distributed-LLM-Edges/
@@ 1 Introduction
Scaling laws [1, 2] have been fundamental to the remarkable success of foundation models, demonstrating a predictable relationship between performance and the expansion of model parameters and training data. These laws have guided the development of increasingly powerful models, from BERT [3] to GPT-4 [4], showing that performance improvements can be achieved through systematic scaling of both model size and training data [5, 6]. However, the continued application of these scaling laws requires ever-increasing amounts of data and computational resources, pushing the boundaries of what is currently feasible [7].
@@
Public data has been the primary fuel driving AI development forward. This field has witnessed an exponential growth in data requirements, from the early success of MNIST [8] with its 70,000 handwritten digits to ImageNet’s revolutionary impact with 14 million labeled images [9]. This trajectory has continued with modern large language models (LLMs) like GPT [4], LLaMA [10], and DeepSeek [11] series, which are trained on trillions of tokens. Recent evidence that LLaMA 3.1’s smallest model (8B) [12] trained on 15 trillion tokens, outperforms LLaMA 2’s largest model (70B) [13] trained on 2 trillion tokens (despite being 10×10$\times$10 × smaller in model size, the 7×7$\times$7 × increase in training data leads to superior performance), demonstrates the paramount importance of data scaling [14, 15]. However, we are witnessing a concerning trend of data exhaustion, where high-quality public data sources are becoming exhausted [16, 17]. [18] argues that human-generated public text data cannot sustain scaling beyond this decade. While recent efforts advocate for training larger models with synthetic data [19], AI-generated content may fail to yield performance improvements [20], also risks polluting public data sources [21]. Moreover, stricter data privacy regulations like GDPR [22] have made data collection increasingly difficult and expensive. This looming data scarcity suggests that scaling laws may hit a wall [23], potentially impeding further AI advancement.
@@
Computational resources has been the primary engine powering AI development. Throughout AI history, major breakthroughs have been closely tied to advances in computing power, from early models requiring single CPUs (with peak performance of 1-2 GFLOPS) to modern GPU clusters. The computational demands have grown exponentially - from BERT-Large’s training requiring 64 TPU v3 chips (providing 420 TFLOPS) [3] to GPT-3’s training on 10,000 V100 GPUs (reaching 28,000 TFLOPS) [5], while training GPT-4 reportedly required over 25,000 NVIDIA A100 GPUs (delivering a staggering 400,000 TFLOPS) [4]. More recent models like Grok 3 push these requirements even further [24]. However, we are approaching physical limits in single-chip performance as Moore’s Law slows down [25]. While massive computing clusters can compensate for individual chip limitations, maintaining such infrastructure incurs astronomical costs - estimated at over \$100M for training GPT-4 [26] - and poses significant environmental concerns due to their enormous energy consumption, with each training run emitting as much CO2 as 500 cars driven for a year [27]. Moreover, this level of computing power has become concentrated among a few tech giants, creating a monopolistic landscape that effectively excludes smaller companies and academic institutions from participating in foundational AI research [28]. This centralization of computing resources presents a significant barrier to innovation and democratization in AI development [29].
@@
In this paper, we propose that leveraging massive distributed edge devices offers a promising solution to overcome both data and computing barriers in AI development. Our analysis (using smartphone as an example) reveals two compelling opportunities: First, edge data generated from smartphones for past 5 years are projected to reach 33.1 EB, offering fresh, diverse, and contextually rich training samples. Second, the collective computing power of edge devices - with smartphones delivering 9,278 EFLOPS for past 5 years - demonstrates the feasibility of distributed model training, as training state-of-the-art models like DeepSeek-v3 would require only about 60,723 users with edge devices working (ideally) in parallel to match its current training setup. Based on these insights, we argue that leveraging these massive distributed edge devices can break barriers of data and computing wall, and everyone can participate in training large models with small edge devices. To support this position, we first analyze the critical challenges of large language models, examining both data bottlenecks (§2.1) and computational monopolization (§2.2). We then explore the hidden potential of massive edge devices, investigating their vast untapped distributed data resources (§3.1) and computational capabilities (§3.2). Building on these insights, we investigate technical approaches for overcoming large model challenges through distributed computing architectures (§4): small language models at edges (§4.1), collaborative inference (§4.2), and collaborative training (§4.4). We then identify two critical open challenges: heterogeneous device model fusion and heterogeneous device compute sharing (§5). Finally, we discuss the societal impact like AI democratization, incentive mechanisms, and environmental benefits of this paradigm shift (§A).
@@ 2.1 The Ceiling of Public Data
The rapid advancement of large language models has created an insatiable appetite for training data. Scaling laws establish that model performance improves predictably with data quantity—a relationship that demands exponentially growing datasets [2]. A canonical example is GPT-3, trained on 300 billion tokens spanning books, web content, and programming code [5]. Current projections suggest dataset sizes grow at 0.38 orders of magnitude (2.4$\times$) annually [30], implying models will require three orders of magnitude more data within a decade.
Despite the internet’s vast textual resources, the total stock of high-quality human-generated text remains bounded. Recent estimates place this limit at approximately $4\times 10^{14}$ tokens [30]. [31] argues that current consumption patterns suggest exhaustion of public text data by 2028, potentially accelerated to 2026 through excessive data reuse during training (a practice termed overtraining). Therefore, the finite nature of publicly available human-generated text data is expected to become a major bottleneck for LLM scaling within the next decade. Despite the current large scale of public data, the risk of data exhaustion is rapidly approaching as data demand continues to grow [32].
@@
Faced with the threat of data exhaustion, researchers have proposed various solutions, among which synthetic data generation is considered one of the most promising approaches. By leveraging LLMs to produce their own training data, researchers envision self-sustaining data ecosystems. Early successes in constrained domains like mathematics and code generation, where automated verification ensures quality, demonstrate potential [33]. Recent work [19] demonstrated that diverse synthetic data enhances the performance of LLMs during both pre-training and fine-tuning.
The adoption of synthetic data faces three fundamental challenges. First, model collapse occurs when models iteratively train on their own outputs, causing gradual divergence from original data distributions. This recursive process amplifies biases and reduces output diversity, ultimately degrading model performance across generations [34, 35]. Second, synthetic data quality remains inherently unverifiable in open-domain contexts. While formal domains like mathematics allow algorithmic validation, natural language lacks objective evaluation standards. The absence of ground-truth verification creates self-referential quality assessments, compromising reliability [36, 20]. Finally, synthetic data struggles to replicate human linguistic diversity. Current methods disproportionately replicate dominant language patterns while underrepresenting cultural nuances and low-frequency expressions. This homogeneity limits their utility for training robust general-purpose models [37].
@@ 2.2 The Monopoly of Computing Resources
Figure 1: Trend of Computational Demand for Model Training. (Data source: [38]).
@@
As large-scale AI models like GPT-4 [4], Llama 3 [12], and DeepSeek-V3 [11] surpass the trillion-parameter scale, the global AI landscape faces severe computational efficiency challenges. As shown in Figure 1, since the deep learning revolution in 2010, AI training demands have grown at a super-exponential rate of 3.9$\times$ per year—an acceleration that intensified with the adoption of the transformer architecture as the industry standard [39]. With the advent of the era of large language models in 2022, the demand for computing power has surged even further, reaching an unprecedented growth rate of 12.8$\times$ per year. This marks a transformative shift in AI computation, where the need for computing power is expanding at an unprecedented pace, pushing the limits of existing hardware and infrastructure.
@@
Moore’s Law, which has driven the growth in computing power for decades, is slowing down as we approach the physical limits of silicon-based chip technology [40]. The difficulty in shrinking transistors has led to diminishing returns in computational performance. As a result, the AI industry is relying more on specialized hardware like GPUs, TPUs, and custom chips to meet growing demands. However, this shift has made high-performance hardware even more expensive and exclusive, further intensifying the gap between organizations with the resources to develop advanced AI models and those without.
@@
The rapid expansion of AI model scales and the surge in computational demand are facing dual constraints in global computing infrastructure. On one hand, bottlenecks in advanced semiconductor manufacturing severely limit the expansion rate of AI data centers. The foundry capacity for wafers at 5nm and below—such as those produced by TSMC—has already been fully booked by leading technology companies until 2026 [41]. Moreover, the construction of new wafer fabs involves long lead times and is further constrained by the global supply chain shortages of critical equipment, such as lithography machines. On the other hand, the exponential increase in chip deployment within individual AI clusters is putting immense pressure on the already limited semiconductor manufacturing capacity, pushing the industry toward its production ceiling [42]. These factors have significantly hindered the continuous expansion of computing power, making it increasingly difficult to scale AI infrastructure sustainably.
@@ 3.1 Massive Data from Edges
As discussed in § 2.1, edge data represents a crucial alternative to synthetic data in addressing the challenge of data exhaustion. Edge data refers to the data generated by edge devices at or near the source of data generation, which typically remains private and localized rather than being publicly accessible. Edge devices encompass a wide range of equipment including Internet of Things (IoTs) sensors, smartphones, wearables, industrial controllers, and other smart devices that process data at the network edge. Data generated at edges offers unique advantages in both data volume and data quality.
Figure 2: Global data volume from 2014 to 2025 and IoT device data volume in 2015 and 2025. (Data sources: Global data volume from [43]; IoT device data volume from [44].)
Figure 3: Smartphone data volume with edge computing market size (right) from 2018 to 2028. (Data sources: Edge computing market [45]; Smartphone data volume from [46].)
@@
According to the statistical data from [43, 44] (as illustrated in Figure 3), the global data volume is projected to reach 182 ZB by 2025 [43], where the data generated by IoT devices is anticipated to increase from 13.6 ZB in 2019 to 79.4 ZB in 2025 [44], elevating its share of the global data volume from 33.2% to 43.6%, showing a particularly pronounced growth in edge-generated data. Over the period from 2015 to 2025, the global data volume exhibited a compound annual growth rate (CAGR) of 35.1%, resulting in an overall increase of 1074.2% and a cumulative total of 871.5 ZB. IoT device data experienced a growth of 483.8% from 2019 to 2025. This trend underscores the increasingly central role of edge-generated data in the global data ecosystem. Beyond IoT devices, smartphones, as a critical source of edge-side data, are also contributing to the steady rise in data volume. As depicted in Figure 3, the estimated smartphone data volume is projected to grow from 5 EB in 2018 to 8 EB by 2028  [^1] These numbers are estimated based on an average user data generation of 1 GB per device. For detailed estimation methodology, refer to Appendix C.. This exponential growth is closely aligned with the rapid expansion of the edge computing market, which is forecasted to surge from \$5.5 billion in 2019 to \$87.9 billion by 2028, representing a remarkable growth rate of 3215.1%. The burgeoning edge computing market has further catalyzed the generation and processing of edge-side data, reinforcing its significance in the broader data landscape 222[47, 48] provide a more comprehensive overview of the global data volume, but we cannot access the statistics data. We appreciate any suggestions for better statistical data sources. .
Edge-generated data has distinctive advantages.
@@
Beyond its impressive quantity, edge data possesses several distinctive characteristics that make it particularly valuable for model training. First, edge data provides enhanced privacy characteristics. Since edge data typically remains local to devices and does not need to be centrally stored, it allows for more privacy-preserving approaches to data utilization. This local-first nature enables compliance with increasingly strict data privacy regulations while still allowing the data to contribute to model training. Second, edge data exhibits superior diversity across multiple dimensions. It encompasses a wide variety of data types from IoT devices, mobile interactions, and personal devices, covering different domains, languages, and user behaviors. This natural diversity provides richer training signals compared to curated public datasets [49]. Third, edge data demonstrates strong real-time capability. Unlike public datasets, which are often updated infrequently, edge devices continuously generate fresh data with low latency [50], offering more up-to-date and relevant training samples. Despite these advantages, edge-generated data can present challenges such as low signal-to-noise ratios, unclean, or potentially harmful content. However, recent advancements in data quality assessment methods (e.g., data prospector [51]), have emerged to identify and filter high-quality data from edge sources, ensuring that only the most reliable and valuable data is selected for model training.
@@
In conclusion, edge data with its explosive growth and distinctive characteristics, is a valuable resource for model pretraining. Its diversity, real-time nature, personalization, and rich context make it an ideal foundation for developing robust and adaptable large-scale models, enhancing their ability to serve real-world applications effectively.
Insight: The smartphone data volume of the past 5 years (before 2025) is projected to reach approximately 33.1 EB, with unique advantages in privacy, diversity, and real-time context, demonstrating the massive data potential of edge for AI model training.